# Robust Hallucination Detection in LLMs via Adaptive Token Selection

### v2

@@ 5.1 Main Results
In Table 1, we evaluate our proposed HaMI by comparing it with eight SotA competing detection methods on four diverse QA datasets in LLMs from two different families of various sizes. As depicted in the table, our approach achieves superior performance compared with other methods in all three LLMs. In particular, HaMI consistently and substantially outperforms all methods that do not require assistance from external LLMs, including Perplexity, CCS, SAPLMA and HaloScope. When compared to approaches that leverage LLMs at the post-generation stage, HaMI achieves similar superiority across all cases except on the SQuAD dataset with LLaMA-3.1-8B, where its performance is comparable to the best score achieved by the SE method.
@@
Comparing the performance on LLaMA models with various scales, we observe that HaMI achieves a notably larger margin of improvement compared to baselines for larger LLaMA-3.13-Instruct-70B. For instance, on the LLaMA-3.1-8B model, HaMI surpasses SE and SAPLMA by an average of 7.4% and 5.5%, respectively. On the larger model, these improvements increase to 11.5% and 8.7%, representing a relative improvement of over 50%.
@@
p(True) kadavath2022language 0.666 0.614 0.650 0.673 0.862 0.749 0.772 0.749 0.563 0.596 0.530 0.573
Perplexity ren2023outofdistribution 0.732 0.649 0.659 0.709 0.720 0.662 0.646 0.675 0.671 0.626 0.619 0.583
SE farquhar2024detecting 0.828 0.787 0.773 0.757 0.795 0.752 0.813 0.800 0.819 0.643 0.769 0.772
MARS bakman-etal-2024-mars 0.766 0.663 0.661 0.706 0.765 0.693 0.700 0.710 0.704 0.648 0.635 0.610
MARS-SE bakman-etal-2024-mars 0.824 0.780 0.777 0.744 0.797 0.743 0.786 0.798 0.794 0.676 0.772 0.769
CCS burns2022discovering 0.675 0.596 0.628 0.662 0.551 0.579 0.615 0.553 0.597 0.575 0.592 0.562
SAPLMA azaria-mitchell-2023-internal 0.835 0.769 0.781 0.821 0.860 0.797 0.836 0.859 0.842 0.672 0.817 0.748
HaloScope du2024haloscope 0.610 0.669 0.637 0.635 0.653 0.621 0.568 0.676 0.656 0.655 0.622 0.556
@@ 5.2 Cross-dataset Generalisation Ability
The ability to generalise across datasets is essential to facilitating real-world applications of LLMs in diverse domains. We conduct experiments on the aforementioned four datasets to assess whether the proposed HaMI can effectively generalise among various datasets, with three strong competing methods as baselines. For each dataset, we report the average AUROC scores of detectors trained on one of the other three datasets. The results are illustrated in Figure 3(a). It is clear that our method HaMI achieves consistently the best generalisation performance across all four datasets, outperforming the best competing method SAPLMA by 7% - 9%. Compared to the within-dataset performance in Table 1, the maximum performance decline in HaMI is no more than $4.5\%$, observed on the BioASQ dataset, which is significantly lower than that of the competing methods. These findings affirm the effectiveness of HaMI as a reliable hallucination detector in unseen datasets.
@@ 5.3 Ablation Study
Furthermore, our ATS module exhibits substantially better robustness to the length of generations than the competing methods. In particular, the average number of generated tokens for the Trivia QA dataset with the LLaMA model is $10$, which is smaller than the other three cases (Trivia QA with Mistral model - $13.5$, SQuAD with LLaMA model - $17.5$, SQuAD with Mistral model - $15$). Despite this variation, the ATS module performs robustly on all four cases and even performs more effectively as the length increases. In contrast, the performance of the predefined token location methods performs unstably and tends to degrade as the length of the generation grows. Moreover, the Last token achieves better performance across all scenarios, and the First token is just the opposite. It suggests that the last token can capture full semantics under specific length conditions, while the first token may miss content in the subsequent predictions.
Figure 4 presents an illustrative example of the scoring results over tokens in a positive bag and a negative bag, where we can observe that there is significant distinguishability between the positive and negative tokens since the maximum scores of the instances in the positive bag is substantially greater than that in the negative bag. Tokens denoted by blue characters have scores that are very close to the largest ones and we find that these tokens can be adaptively identified by our method as concentrated answers in comparison to ground-truth answers. Additionally, it is noted that while selected tokens are associated with the exact answer, they may appear at any location in their vicinity, which can be captured by our smoothness loss.
@@
Here we systematically examine three uncertainty measuring methods presented in Section 4.2 for the internal representation enhancement as detailed in Eq. 7, 8 and 9 respectively. We use a simple combination strategy as defined in Eq. 10. In the right panel of Table 2, we can observe that both $P_{\text{uncertainty}}^{s}$ and $P_{\text{uncertainty}}^{c}$ contribute to improved detection capabilities over the baseline (i.e.,, the original feature representations) while the token-level uncertainty $P_{\text{uncertainty}}^{t}$ yields limited effectiveness. Specifically, the enhancements attributed to semantic consistency $P_{\text{uncertainty}}^{c}$ are the most significant, exhibiting improvements up to $6.7\%$. These observations suggest that the effectiveness of uncertainty-based enhancement is tied to the inherent ability of the uncertainty metric to distinguish hallucinations from correct responses. Notably, although the improvements observed with $P_{\text{uncertainty}}^{s}$ are less pronounced, it surpasses the performance of SotA multi-sampling approaches (such as SE in Table 1), without incurring the costs associated with multiple generations and external LLM employments. This highlights the potential of HaMI for deployments in various practical environments that involve external tools or not.
@@
Figure 5(b) presents results based on various choice of $\lambda$ used in Eq. 10. It is clear that the model’s performance is steadily improved as $\lambda$ changed from 0.0 towards 1.0. Beyond this point, performance becomes stable.
@@ A.1 Setup
For the main results, our hallucination detector $f_{\theta}(\cdot)$ utilises a two-layer MLP with a hidden dimension of 256. The first linear layer is followed by a BatchNorm and ReLU activations and the second linear layer is accompanied by a sigmoid output. As shown in Figure 3b in our paper, our method is relatively insensitive to variations in the number of layers and the hidden dimension size. At the training stage, we use Adam optimiser. For the input representations of our detector, we utilise representations extracted from a single layer and the layer index is determined by the results of the validation set.
We implement our method using PyTorch 2.6.0 [36] and transformers 4.51.3 [43] and conduct all experiments on NVIDIA A100 GPUs. Based on experimental records, the estimated total training and inference time for each dataset under the main experimental settings is as follows: 0.40 GPU-hours for LLaMA-3.1-8B, 0.45 GPU-hours for Mistral-Nemo-Instruct (12B) and 0.65 GPU-hours for LLaMA-3.13-Instruct-70B with 8-bit quantisation. It is important to note that all these results are based on the assumption that the semantic uncertainty score $P_{\text{uncertainty}}^{c}$ is readily available. Otherwise, computing this score involves generating multiple responses and evaluating entailment across them, which introduces nontrivial latency, approximately 7.6 seconds per question for LLaMA-3.1-8B, 8.6 seconds for Mistral-Nemo-Instruct (12B), and 46.1 seconds for LLaMA-3.13-Instruct-70B.
@@ A.2 Prompts for generation and evaluation
Appendix B Study on $k$ for the top $k$ highest hallucination score selection
As presented in Section 4.1 in our paper, we select token instances with the top $k$ highest hallucination scores as the salient tokens in each sequence. $k$ is dynamically determined by the generated token length $l$ as follows:
@@ Appendix B Study on $k$ for the top $k$ highest hallucination score selection
where $r_{k}$ is set to 0.1 by default. In this section, we investigate the sensitivity of our method to the choice of $r_{k}$ in the adaptive token selection process. As shown in Figure 8, HaMI exhibits stable performance across a set of $r_{k}$, ranging from 0.00 to 0.20, on all four datasets with LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B). Notably, smaller $r_{k}$ values already yield strong results, which implies that only a few high-salience tokens are sufficient for effective hallucination discrimination.
Figure 8: Study on different choice of $r_{k}$ across four datasets with two LLMs.
Appendix C Limitations and Broader Impacts

### v3

@@ 5.1 Main Results
In Table 1, we evaluate our proposed HaMI by comparing it with eight SotA competing detection methods on four diverse QA datasets in LLMs from two different families of various sizes. As depicted in the table, our approach achieves superior performance compared with other methods in all three LLMs. In particular, HaMI consistently and substantially outperforms all methods that do not require assistance from external LLMs, including Perplexity, CCS, SAPLMA, HaloScope, CED, and LLM-Check. When compared to approaches that leverage LLMs at the post-generation stage, HaMI achieves similar superiority across all cases except on the SQuAD dataset with LLaMA-3.1-8B, where its performance is comparable to the best score achieved by the SE method.
@@
Comparing the performance on LLaMA models with various scales, we observe that HaMI achieves a notably larger margin of improvement compared to baselines for larger LLaMA-3.3-Instruct-70B. For instance, on the LLaMA-3.1-8B model, HaMI surpasses SE and SAPLMA by an average of 7.4% and 5.5%, respectively. On the larger model, these improvements increase to 11.5% and 8.7%, representing a relative improvement of over 50%.
@@
p(True) [25] 0.666 0.614 0.650 0.673 0.862 0.749 0.772 0.749 0.563 0.596 0.530 0.573
Perplexity [42] 0.732 0.649 0.659 0.709 0.720 0.662 0.646 0.675 0.671 0.626 0.619 0.583
SE [17] 0.828 0.787 0.773 0.757 0.795 0.752 0.813 0.800 0.819 0.643 0.769 0.772
MARS [6] 0.766 0.663 0.661 0.706 0.765 0.693 0.700 0.710 0.704 0.648 0.635 0.610
MARS-SE [6] 0.824 0.780 0.777 0.744 0.797 0.743 0.786 0.798 0.794 0.676 0.772 0.769
CCS [9] 0.675 0.596 0.628 0.662 0.551 0.579 0.615 0.553 0.597 0.575 0.592 0.562
SAPLMA [5] 0.835 0.769 0.781 0.821 0.860 0.797 0.836 0.859 0.842 0.672 0.817 0.748
HaloScope [14] 0.610 0.669 0.637 0.635 0.653 0.621 0.568 0.676 0.656 0.655 0.622 0.556
CED [29] 0.745 0.776 0.710 0.695 0.693 0.680 0.669 0.674 0.709 0.699 0.714 0.703
LLM-Check-h [45] 0.666 0.614 0.610 0.673 0.696 0.659 0.673 0.668 0.678 0.639 0.642 0.702
LLM-Check-a [45] 0.651 0.644 0.626 0.675 0.715 0.677 0.690 0.670 0.691 0.657 0.653 0.713
@@ 5.2 Cross-dataset Generalisation Ability
The ability to generalise across datasets is essential to facilitating real-world applications of LLMs in diverse domains. We conduct experiments on the aforementioned four datasets to assess whether the proposed HaMI can effectively generalise among various datasets, with three strong competing methods in the line of internal-representation analysis as baselines. For each dataset, we report the average AUROC scores of detectors trained on one of the other three datasets. The results are illustrated in Figure 3(a). It is clear that our method HaMI consistently achieves the best generalisation performance across all four datasets, outperforming the best competing method SAPLMA by 7% - 9%. Compared to the within-dataset performance in Table 1, the maximum performance decline in HaMI is no more than $4.5\%$, observed on the BioASQ dataset, which is significantly lower than that of the competing methods. We also compare the generalisation capability of HaMI with training-free baselines. The results for HaMI trained on one dataset and evaluated on others separately are presented in Table 2, where the results for the training-free baselines are the best AUROC of training-free methods—p(True), Perplexity, SE, MARS, MARS-SE, CED, LLM-Check-h, and LLM-Check-a—on each dataset. Notably, across all settings, HaMI can consistently
Trivia QA SQuAD NQ BioASQ
Trained on Trivia QA - 0.802 0.818 0.796
Trained on SQuAD 0.850 - 0.794 0.804
Trained on NQ 0.870 0.804 - 0.791
Trained on BioASQ 0.853 0.790 0.781 -
Training-Free Baselines 0.828 0.787 0.777 0.757
Table 2: Cross-dataset generalization capability compared with training-free baselines.
outperform training-free/non-learnable baselines on unseen target datasets, regardless of which dataset it is trained on. The observation is similar to the one with the results aggregated over multiple detectors. These findings affirm the effectiveness of HaMI as a reliable hallucination detector in unseen datasets.
@@ 5.3 Ablation Study
Furthermore, our ATS module exhibits substantially better robustness to the length of generations than the competing methods. In particular, the average number of generated tokens for the Trivia QA dataset with the LLaMA model is $10$, which is smaller than the other three cases (Trivia QA with Mistral model - $13.5$, SQuAD with LLaMA model - $17.5$, SQuAD with Mistral model - $15$). Despite this variation, the ATS module performs robustly on all four cases and even performs more effectively as the length increases. In contrast, the performance of the predefined token location methods is unstable and tends to degrade as the length of the generation grows. Moreover, the Last token achieves better performance across all scenarios, and the First token is just the opposite. It suggests that the last token can capture full semantics under specific length conditions, while the first token may miss content in the subsequent predictions.
To further support the validity of the token selection process, we conducted an additional study, a manual evaluation using 100 randomly sampled examples from the Trivia QA dataset. In this analysis, we manually annotated hallucinated tokens and measured the token detection recall rate of our method throughout training. Specifically, HaMI can achieve a recall rate of 0.84 as training progresses. This suggests that our approach enables the model to identify salient tokens even without token-level supervision. Figure 4 presents an illustrative example of the scoring results over tokens in a positive bag and a negative bag, where we can observe that there is significant distinguishability between the positive and negative tokens since the maximum scores of the instances in the positive bag are substantially greater than those in the negative bag. Tokens denoted by blue characters have scores that are very close to the largest ones and we find that these tokens can be adaptively identified by our method as concentrated answers in comparison to ground-truth answers. Additionally, it is noted that while selected tokens are associated with the exact answer, they may appear at any location in their vicinity, which can be captured by our smoothness loss.
@@
Here we systematically examine three uncertainty measuring methods presented in Section 4.2 for the internal representation enhancement as detailed in Eq. 7, 8 and 9 respectively. We use a simple combination strategy as defined in Eq. 10. In the right panel of Table 3, we can observe that both $P_{\text{uncertainty}}^{s}$ and $P_{\text{uncertainty}}^{c}$ contribute to improved detection capabilities over the baseline (i.e., the original feature representations) while the token-level uncertainty $P_{\text{uncertainty}}^{t}$ yields limited effectiveness. Specifically, the enhancements attributed to semantic consistency $P_{\text{uncertainty}}^{c}$ are the most significant, exhibiting improvements up to $6.7\%$. These observations suggest that the effectiveness of uncertainty-based enhancement is tied to the inherent ability of the uncertainty metric to distinguish hallucinations from correct responses. Notably, although the improvements observed with $P_{\text{uncertainty}}^{s}$ are less pronounced, it surpasses the performance of SotA multi-sampling approaches (such as SE in Table 1), without incurring the costs associated with multiple generations and external LLM employments. This highlights the potential of HaMI for deployments in various practical environments that involve external tools or not.
@@
Figure 5(b) presents results based on various choice of $\lambda$ used in Eq. 10. It is clear that the model’s performance is steadily improved as $\lambda$ changes from 0.0 towards 1.0. Beyond this point, performance becomes stable.
@@ 6 Conclusion
Acknowledgments
This research is supported by A*STAR under its MTC YIRG Grant (M24N8c0103), the Ministry of Education, Singapore under its Tier-1 Academic Research Fund (24-SIS-SMU-008), and the Lee Kong Chian Fellowship.
@@ A.1 Setup
For the main results, our hallucination detector $f_{\theta}(\cdot)$ utilizes a two-layer MLP with a hidden dimension of 256. The first linear layer is followed by a BatchNorm and ReLU activations and the second linear layer is accompanied by a sigmoid output. As shown in Figure 3b in our paper, our method is relatively insensitive to variations in the number of layers and the hidden dimension size. At the training stage, we use Adam optimiser. For the input representations of our detector, we utilise representations extracted from a single layer and the layer index is determined by the results of the validation set.
We implement our method using PyTorch 2.6.0 [39] and transformers 4.51.3 [48] and conduct all experiments on NVIDIA A100 GPUs. Based on experimental records, the estimated total training and inference time for each dataset under the main experimental settings is as follows: 0.40 GPU-hours for LLaMA-3.1-8B, 0.45 GPU-hours for Mistral-Nemo-Instruct (12B) and 0.65 GPU-hours for LLaMA-3.3-Instruct-70B with 8-bit quantisation. It is important to note that all these results are based on the assumption that the semantic uncertainty score $P_{\text{uncertainty}}^{c}$ is readily available. Otherwise, computing this score involves generating multiple responses and evaluating entailment across them, which introduces nontrivial latency, approximately 7.6 seconds per question for LLaMA-3.1-8B, 8.6 seconds for Mistral-Nemo-Instruct (12B), and 46.1 seconds for LLaMA-3.3-Instruct-70B. In this work, we also propose a lightweight variant, HaMI, using only a single generation, to eliminate the influence on computational efficiency while maintaining superior performance.
@@ A.2 Prompts for Generation and Evaluation
Appendix B Study on the Effectiveness of the Smoothness Loss
The smoothness loss is designed under the assumption of continuity in hallucination scores between adjacent tokens. We investigate the effectiveness of smoothness loss on four benchmarks with LLaMA-3.1-8B model. Main results on each individual dataset and cross-dataset results are respectively presented in Table 5 and Table 5 below. The results are averaged over three independent runs. As evidenced in the tables, incorporating the smoothness loss consistently yields improved performance across all the cases. On average, the smoothness loss consistently enhances HaMI, increasing the performance by $0.7\%$ and $1.1\%$ on the within-dataset and cross-dataset scenarios, respectively. This demonstrates that the contribution of the smoothness loss is more pronounced in the more challenging cross-dataset setting. These consistent improvements highlight that the smoothness loss plays a rather positive role in helping achieve a more robust and generalizable HaMI, which is particularly valuable in cross-domain scenarios.
$\mathcal{L}_{MIL}$ $\mathcal{L}_{smooth}$ Trivia QA SQuAD NQ BioASQ
[…]
Table 4: Ablation study on the MIL and smoothness losses.
$\mathcal{L}_{MIL}$ $\mathcal{L}_{smooth}$ Trivia QA SQuAD NQ BioASQ
[…]
Table 5: Ablation study on the MIL and smooth losses under the cross-dataset setting.
Appendix C Study on Top-$k$ Token Selection
HaMI employs a hard top-$k$ selection strategy in its MIL formulation that functions similarly to max pooling during training. This design aligns with the core assumption in multi-instance learning (MIL) that if there are any positive instances, the bag should be labelled as positive. This principle is particularly relevant to our detection task, where the salient hallucinated tokens appear sparsely within generated sequences. While there exist more sophisticated and learnable aggregation techniques, previous research has indicated that, given a sufficient number of training bags, different aggregation methods for a bag tend to converge to similar performance [20]. Based on these observations and the remarkable performance of our method in our experimental results, despite being simple, the hard top-$k$ strategy offers a direct, yet effective solution for the detection task. As presented in Section 4.1, $k$ is dynamically determined by the generated token length $l$ as follows:
@@ Appendix C Study on Top-$k$ Token Selection
where $r_{k}$ is set to 0.1 by default. In this section, we investigate the sensitivity of our method to the choice of $r_{k}$ in the adaptive token selection process. As shown in Figure 8, HaMI exhibits stable performance across a set of $r_{k}$, ranging from zero to 0.20, on all four datasets with LLaMA-3.1-8B and Mistral-Nemo-Instruct (12B). HaMI can perform stably across different choices of $r_{k}$. Notably, smaller $r_{k}$ values already yield strong results, which implies that only a few high-salience tokens are sufficient for effective hallucination discrimination.
Figure 8: Study on different choices of $r_{k}$ across four datasets with two LLMs.
Appendix D Evaluation results with BLEURT score
In addition to labels identified by GPT-4.1 [1], we also use the BLEURT-threshold [43] method to generate labels and comparing HaMI with five representative baselines. Results are reported in Table 6, which shows that our methods HaMI and HaMI can also consistently outperform the baselines using the BLEURT-based labelled datasets. These results provide empirical evidence for the robustness of HaMI from another perspective (i.e., a data labelling perspective).
Method Trivia QA SQuAD NQ BioASQ
Perplexity 0.663 0.638 0.586 0.662
SE 0.709 0.813 0.704 0.714
CCS 0.682 0.657 0.738 0.683
SAPLMA 0.897 0.898 0.867 0.873
HaloScope 0.792 0.845 0.852 0.830
HaMI 0.900 0.928 0.909 0.918
HaMI 0.909 0.937 0.920 0.922
Table 6: AUROC results with labels generated by the BLEURT method. The results are based on LLaMA-3.1-8B.
Appendix E Limitations and Broader Impacts